# Scattering Cross Section Formula Derived From Macroscopic Model of Detectors

### v1

@@ 6.2 $N$-Particle Systems
As shown in [13], the density of $\mathbb{P}_{WOD}$ in the far-field regime can also be expressed in terms of a multi-time wave function [21], [39, sec. 7.4.1]
@@ 6.3.3 Asymptotic Trajectories
While in the non-relativistic case [32] and in the Dirac case with wave functions from the positive-energy subspace of the free Dirac Hamiltonian, the Bohmian world lines of free particles asymptotically become straight lines (i.e., the velocity becomes constant), this is not so for general solutions of the free Dirac equation, as we derive in Appendix B. Rather, the Bohmian world lines asymptotically become elliptical helices whose axes are timelike straight lines in direction $k^{\mu}$, with period length (the time required for one winding in the rest frame of the axis) given by $\pi\hbar/mc^{2}$ (the Compton wave length divided by $2c$), which for an electron has the value of $4.047\times 10^{-21}\,\text{s}$, and major radius of the ellipse (spatial size of a winding in the rest frame of the axis) no larger than $\hbar/2mc$ (the Compton wave length divided by $4\pi$), which for an electron is $1.93\times 10^{-13}$ m. Holland [18, Sec. 12.3.3] suggested in a similar situation to regard such a periodic motion around a straight line as a Bohmian version of Schrödinger’s [33] concept of Zitterbewegung.333“Zitterbewegung” (trembling motion) is a name for the fact that the Heisenberg-evolved position operator for a single Dirac particle has an oscillatory time dependence with the Compton frequency (see, e.g., [35, Sec. 1.6.1]). It seems that Holland was not aware of the general fact that asymptotically as $t\to\infty$ for the free Dirac equation, the trajectories keep oscillating forever (i.e., become helices), a fact that arises from the fact that the Fourier mode with wave vector $k^{\mu}$ and the one with $k^{\mu}$ will never separate in space but will keep jointly influencing the particle’s motion. Instead, his example consists of a superposition of two plane waves, one with wave vector $(k^{0},\boldsymbol{k})$ with $k^{0}=(\boldsymbol{k}^{2}+m^{2}c^{2}/\hbar^{2})^{1/2}$ and the other with $(-k^{0},\boldsymbol{k})$, which are two Fourier modes that will ultimately get separated in space for every normalizable wave function, as they move at different velocities $c\boldsymbol{k}/k^{0}\neq-c\boldsymbol{k}/k^{0}$. As soon as they get separated, the oscillatory motion stops and gets replaced by a straight trajectory; they do not get separated in Holland’s example only because the example is not normalizable.
@@ Appendix C Reflection Coefficient for Imaginary Potential Step in the Dirac Equation
That is, the additional potential applies in the half space $\{x^{3}\geq 0\}$ and is given there by a constant skew-adjoint matrix $i\lambda A_{\mu}\gamma^{0}\gamma^{\mu}=-i\lambda A_{0}\,I+i\lambda\boldsymbol{A}\cdot\boldsymbol{\alpha}$. As in Section 4, the key step is to find the eigenfunctions of the operator (156). To this end, we need the following statement, which is well known [35, (1.42)] for real $\boldsymbol{k}$, also for complex $\boldsymbol{k}$.
@@
Proof. We first use the known facts $\beta^{2}=I$, $\{\alpha_{i},\beta\}=0$, and $\{\alpha_{i},\alpha_{j}\}=2\delta_{ij}I$ [35, (1.6)] to show that
@@
It is known [35, (1.228)] that $\boldsymbol{a}\cdot\boldsymbol{\alpha}$ and $\boldsymbol{b}\cdot\boldsymbol{\alpha}$ commute for real vectors $\boldsymbol{a},\boldsymbol{b}$ if and only if $\boldsymbol{a}\,||\,\boldsymbol{b}$; thus, $M$ is normal (i.e., $MM^{\dagger}=M^{\dagger}M$, or the self-adjoint and skew-adjoint parts of $M$ commute) if and only if $\mathrm{Re}\,\boldsymbol{k}\,||\,\mathrm{Im}\,\boldsymbol{k}$. On the other hand, a diagonalizable matrix $M$ is normal if and only if all eigenspaces are mutually orthogonal. $\square$

### v2

@@ 6.2 $N$-Particle Systems
As shown in [14], the density of $\mathbb{P}_{WOD}$ in the far-field regime can also be expressed in terms of a multi-time wave function [23], [42, sec. 7.4.1]
@@ 6.3.3 Asymptotic Trajectories
While in the non-relativistic case [34] and in the Dirac case with wave functions from the positive-energy subspace of the free Dirac Hamiltonian, the Bohmian world lines of free particles asymptotically become straight lines (i.e., the velocity becomes constant), this is not so for general solutions of the free Dirac equation, as we derive in Appendix B. Rather, the Bohmian world lines asymptotically become elliptical helices whose axes are timelike straight lines in direction $k^{\mu}$, with period length (the time required for one winding in the rest frame of the axis) given by $\pi\hbar/mc^{2}$ (the Compton wave length divided by $2c$), which for an electron has the value of $4.047\times 10^{-21}\,\text{s}$, and major radius of the ellipse (spatial size of a winding in the rest frame of the axis) no larger than $\hbar/2mc$ (the Compton wave length divided by $4\pi$), which for an electron is $1.93\times 10^{-13}$ m. Holland [20, Sec. 12.3.3] suggested in a similar situation to regard such a periodic motion around a straight line as a Bohmian version of Schrödinger’s [35] concept of Zitterbewegung.444“Zitterbewegung” (trembling motion) is a name for the fact that the Heisenberg-evolved position operator for a single Dirac particle has an oscillatory time dependence with the Compton frequency (see, e.g., [37, Sec. 1.6.1]). It seems that Holland was not aware of the general fact that asymptotically as $t\to\infty$ for the free Dirac equation, the trajectories keep oscillating forever (i.e., become helices), a fact that arises from the fact that the Fourier mode with wave vector $k^{\mu}$ and the one with $k^{\mu}$ will never separate in space but will keep jointly influencing the particle’s motion. Instead, his example consists of a superposition of two plane waves, one with wave vector $(k^{0},\boldsymbol{k})$ with $k^{0}=(\boldsymbol{k}^{2}+m^{2}c^{2}/\hbar^{2})^{1/2}$ and the other with $(-k^{0},\boldsymbol{k})$, which are two Fourier modes that will ultimately get separated in space for every normalizable wave function, as they move at different velocities $c\boldsymbol{k}/k^{0}\neq-c\boldsymbol{k}/k^{0}$. As soon as they get separated, the oscillatory motion stops and gets replaced by a straight trajectory; they do not get separated in Holland’s example only because the example is not normalizable.
@@ Appendix C Reflection Coefficient for Imaginary Potential Step in the Dirac Equation
That is, the additional potential applies in the half space $\{x^{3}\geq 0\}$ and is given there by a constant skew-adjoint matrix $i\lambda A_{\mu}\gamma^{0}\gamma^{\mu}=-i\lambda A_{0}\,I+i\lambda\boldsymbol{A}\cdot\boldsymbol{\alpha}$. As in Section 4, the key step is to find the eigenfunctions of the operator (170). To this end, we need the following statement, which is well known [37, (1.42)] for real $\boldsymbol{k}$, also for complex $\boldsymbol{k}$.
@@
Proof. We first use the known facts $\beta^{2}=I$, $\{\alpha_{i},\beta\}=0$, and $\{\alpha_{i},\alpha_{j}\}=2\delta_{ij}I$ [37, (1.6)] to show that
@@
It is known [37, (1.228)] that $\boldsymbol{a}\cdot\boldsymbol{\alpha}$ and $\boldsymbol{b}\cdot\boldsymbol{\alpha}$ commute for real vectors $\boldsymbol{a},\boldsymbol{b}$ if and only if $\boldsymbol{a}\,||\,\boldsymbol{b}$; thus, $M$ is normal (i.e., $MM^{\dagger}=M^{\dagger}M$, or the self-adjoint and skew-adjoint parts of $M$ commute) if and only if $\mathrm{Re}\,\boldsymbol{k}\,||\,\mathrm{Im}\,\boldsymbol{k}$. On the other hand, a diagonalizable matrix $M$ is normal if and only if all eigenspaces are mutually orthogonal. $\square$
@@
Funding. R.K. gratefully acknowledges financial support from German Academic Exchange Service (DAAD) (funding programme/ID: Research Grants - Doctoral Programmes in Germany, 2024/25 (57693453)).
Acknowledgments. We thank Lawrence Frolov, Fabian Nolte, and Shadi Tahvildar-Zadeh for helpful discussion.